OpenStudy (anonymous):

Find the indicated arc length. The arc corresponding to a central angle of π/3 in a circle of radius 6 inches

jimthompson5910 (jim_thompson5910):

The circumference of this circle is C = 2*π*r C = 2*π*6 C = 12π The angle π/3 is (π/3)/(2π) = 1/6 of the way to completing a full circle, so you multiply the circumference with 1/6 to get 12π*(1/6) = 12π/6 = 2π So the arc is 2π units long

OpenStudy (anonymous):

thanks so much :) i really appreciate it!! so 6.28 degrees is the length of the arc?

jimthompson5910 (jim_thompson5910):

no 6.28 units (not degrees or radians)

jimthompson5910 (jim_thompson5910):

oh in this case, the units are inches so 6.28 inches roughly

OpenStudy (anonymous):

oh ok i see, thanks so much

OpenStudy (anonymous):

so 6.28 is the answer? or is there more steps afterwards

jimthompson5910 (jim_thompson5910):

degrees and radians apply to angles and not lengths

jimthompson5910 (jim_thompson5910):

no that's pretty much it, 6.28 is a rough approximation

OpenStudy (anonymous):

thank you so much for the explanation i really appreciate it!! :)

jimthompson5910 (jim_thompson5910):

you're welcome